OpenStudy (anonymous):

\sqrt{x^{2}+3}+\sqrt{x+3}=5

OpenStudy (mathmath333):

\(\sqrt{x^{2}+3}+\sqrt{x+3}=5\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but I need doing post.Can you help me

OpenStudy (aum):

Square both sides as a first step.

OpenStudy (anonymous):

yes I square

OpenStudy (aum):

What do you get?

OpenStudy (anonymous):

You can not solve for me?

OpenStudy (anonymous):

So I tried every possible way

OpenStudy (aum):

Can you square the left side?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

but it's 25

OpenStudy (anonymous):

I use application \[2\sqrt{AB}=C-A-B\]

OpenStudy (aum):

It might be easier if we take one radical to the other side: \[ \sqrt{x^{2}+3}+\sqrt{x+3}=5 \\ \sqrt{x^{2}+3} = 5- \sqrt{x+3} \\ \text{Square both sides:} \\ x^2 + 3 = 25 + x + 3 - 10\sqrt{x+3} \\ x^2 + 3 - 25 - x - 3 = - 10\sqrt{x+3} \\ x^2 - x - 25 = - 10\sqrt{x+3} \\ \text{Square both sides again:} \]Can you do that part?

OpenStudy (anonymous):

yes I worked to do so still can not calculate when I put t

OpenStudy (anonymous):

Because if it's a 4 set t hat do I do for the next star?

OpenStudy (anonymous):

?

OpenStudy (aum):

\[ x^2 - x - 25 = - 10\sqrt{x+3} \\ \text{Square both sides again:} \\ x^4 + x^2 + 625 - 2x^3 + 50x - 50x^2 = 100x + 300 \\ x^4 - 2x^3 - 49x^2 - 50x +325 = 0 \]

OpenStudy (anonymous):

Well I was out there, but when they do not get put t

OpenStudy (aum):

Are you allowed to use graphing calculators? If allowed we may not even need these steps and just plot the original equation.

OpenStudy (anonymous):

Yeah but my teacher began to explain and not allowed to use computers or rounding

OpenStudy (aum):

Graphing calculators allowed or not allowed?

OpenStudy (anonymous):

not allowed

OpenStudy (aum):

What chapter is this problem from? Numerical Analysis? Give some background so I get an idea what is allowed here.

OpenStudy (aum):

?????

OpenStudy (anonymous):

Well this is all based solution to find x such that this equation based experience satisfying x = 5, is not allowed to rounding or computer use

OpenStudy (aum):

Yes we have to solve for x. But we have a 4th degree equation. We can find solutions by trial and error. Is that allowed? We can try x = 1, x = 2, x = 3, etc, until we find two values where the function changes sign. Then there must be a root between those two numbers.Then we refine it further and find a solution to the first decimal place , then second decimal place, etc. It will take a lot of time and a lot of trial and error.

OpenStudy (anonymous):

Is it because my math teacher for only 2.11 Acceptance very odd but I began to find I do not know how to prove

OpenStudy (aum):

Oh, wait a minute! They gave you x = 2.11 and asked you to prove that it is a solution for the given equation?

OpenStudy (aum):

Are you asked to prove x = 2.11 is a solution?

OpenStudy (anonymous):

No, I press the computer to try to find the test before, but no way to prove it with this experiment or solve all the equations

OpenStudy (anonymous):

Shift +solve is Answer but I don't know solve problems

OpenStudy (aum):

Solving this problem requires one of the following: !. Graphing Calculator 2. Online graphing calculator 3. Trial and error method If the first two are not allowed I can only think of the third choice to solve this problem.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

I will go to bed now good day for you.Thank you to help me

OpenStudy (aum):

You are welcome.

OpenStudy (aum):

Trial and error method: \[ f(x) = \sqrt{x^{2}+3}+\sqrt{x+3} =5 \\ f(0) = \sqrt{0^{2}+3}+\sqrt{0+3} = 3.46 \\ f(1) = \sqrt{1^{2}+3}+\sqrt{1+3} = 4 \\ f(2) = \sqrt{2^{2}+3}+\sqrt{2+3} = 4.88 \\ \text{Answer getting closer to 5...} \]

OpenStudy (aum):

\[ f(x) = \sqrt{x^{2}+3}+\sqrt{x+3} =5 \\ f(2.1) = \sqrt{2.1^{2}+3}+\sqrt{2.1+3} = 4.98 \\ f(2.11) = \sqrt{2.11^{2}+3}+\sqrt{2.11+3} = 4.99 \\ f(2.12) = \sqrt{2.12^{2}+3}+\sqrt{2.12+3} = 5.0003 \\ x \approx 2.12 \]

OpenStudy (anonymous):

thanks